question 1= f(x)=x/x+1; function domen and range =?
queation 2= f(x)=x-3/2x+1; function domen and range =?

Answer by josgarithmetic(39613)

-----
question 1= f(x)=x/x+1; function domen and range =?
queation 2= f(x)=x-3/2x+1; function domen and range =?
-----


(1)
f(x)=x/(x+1)

Domain is all real numbers less than negative one and greater than negative one; the set of possible x values must EXCLUDE 1.

Range is real numbers less than 1 and greater than 1 and MUST EXCLUDE 1.






-------------------------------------------------------
Do not write like this: question 1= f(x)=x/x+1

Write like this:
Question #1
f(x)=x/(x+1)

notice that you probably, most certainly needed those parentheses. Otherwise what you had written as x/x+1 would mean 1+1 or 2.